OpenStudy (dinmosiaren):

what is the inverse of y=2x+3

OpenStudy (anonymous):

Instead of multiplying by 2 and adding 3, the inverse is to subtract 3 then divide by 2.

OpenStudy (dinmosiaren):

okay so y=-3 +x/2?

OpenStudy (dinmosiaren):

or x/2-3

OpenStudy (dinmosiaren):

y=x/2-3

OpenStudy (anonymous):

\[f^{-1} (x)= \frac{ x-3 }{ 2 }\]

OpenStudy (anonymous):

y=(x-3)/2

OpenStudy (dinmosiaren):

oh okay cool thanks